INAF LBC+MODS+LUCI queue observing, 2012 October 18-24

Observer: M. Faccini, E. Sani, K. Boutsia
Telescope Operator: D. Gonzalez Huerta and G.Bechetti
Support Astronomer: D. Thompson and M. Edwards
Instrument: MODS + LBC + LUCI

October 18-19

Project_RUN Objects Filter ExpTime Notes
Lardo_5B M2 SGB blue_grating 2hr15min seeing variable, ~1.0" for 2 exposures and 1.2"-1.5" for the other, probably because of passing clouds.
Giallongo_27A MS0015 g-sloan, r-sloan 1hr50min after allseeing blue:1.1" / red:1.0" - problems with guiding and collimation, fringing features in some red images
Giallongo_27B MAC0717 g-sloan, r-sloan 1hr40min after allseeing blue:1.0" / red:1.0" - problems with guiding and collimation, fringing features in some red images
Standard bd284211 blue_grating   airmass 1.1

October 19-20

Project_RUN Objects Filter ExpTime Notes
Hunt_8G GRB080325 HKspec, 200H+K 20min we are not sure that the target is within the slit
Sozzeti WASP33 HKspec, 200H+K - we are not able to do the mask acquisition because the field does not seem to match the mask - after 1hr of effort we abort
Turatto_07A SN2009ip HKspec - target too low and we loose guide star in one offset - abort
Cresci_28A DSF2237-A H, 210_zJHK 1hr flat +arc +Tell
La Franca_A1 2MSXJ506+32 zJspec, 200 H+K 1hr flat + arc +Tell - COMPLETED
Hunt_8A GRB020127 HKspec, 200H+K - target too faint to identify in the acq image - we need a blind offset -abort
La Franca_A2 2MSXJ09+45 zJspec, 200H+K 1hr flat + arc +Tell - COMPLETED

October 20-21

Project_RUN Objects Filter ExpTime Notes
Lardo_5B M2 RGB blue_grating 1hr target completed - seeing within specs
Lardo_5B M2 SGB blue_grating 45min repeated 1exp because the first night there were clouds during the second exposure - COMPLETED
Standard bd284211 blue_grating    
Gilli C 2XMMJ22 red_grating 45min we had to close dome for humidity after the 3rd exposure
Standard feige red_grating    

October 21-22

Project_RUN Objects Filter ExpTime Notes
Gilli 18C 2XMMJ22 red-grating 1h 30m we repeated the whole program because the seeing is hight - considered COMPLETE
Standard Feige 110 red-grating    
Fritz 4B GG94 dual-grating 2h 40m we took an extra exposure because of seeing being slightly out of specs - considered COMPLETE
Standard G191B2b dual-grating / red-grating    
Giallongo 27A MS0015 dual channel 15m blue side jumping again - we had to close for humidity

October 22-23

Project_RUN Objects Filter ExpTime Notes
Standard Hipp104320 H/210_zJHK   We repeat the telluric acquisition and integration because the spectrum appear too faint
Cresci 18A DSF2237a H/210_zJHK 1h10m flat+arc+Tell
Sozzetti 22B WASP33b HKspec/200_H+K 6hr While the target is transiting, the flexure compensation cannot handle how fast the rotator is moving - COMPLETE

October 23-24

Project_RUN Objects Filter ExpTime Notes
Hunt_8A GRB020127     keep trying to point to the target but we cannot guide because of clouds - impossible to observe
La Franca A3 NGC3079 zJspec, 200 H+K, 1hr problems with the position angle, flat+arc+Tell - COMPLETE

October 24-25

Project_RUN Objects Filter ExpTime Notes
LaFranca_6 3C403 zJspec, 200_H+K 12m Test to check the slit alignment and the saturation of the reference star
Turatto_7A SN2009ip zJspec, 200_H+K 1hr Check of the slit alignment. No collimation. flat+arc+Tell - COMPLETE
Turatto_7A SN2009ip HKspec, 200_H+K 45m No collimation. No guiding, flat+arc+Tell - COMPLETE
Giallongo_A MS0015 g-sloan, r-sloan 30min out of spec
Giallongo_B MAC0717 g-sloan, r-sloan 2hr after allseeing we have median seeing of 1.2" for 1hr15min of blue / red: 1.0" - COMPLETE
Standard SA98 g-sloan, r-sloan   airmass 1.2
Standard SA101 g-sloan, r-sloan   airmass 1.4 - background too high, not sure it is usable

-- KonstantinaBoutsia - 28 Oct 2012
Topic revision: r4 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback